### v1

# Slow-phonon control of spin Edelstein effect in Rashba $d$-wave altermagnets

Abstract
Altermagnets have zero net magnetization yet feature spin-split bands that host spin-polarized states. Here, we investigate how slow lattice vibrations (phonons) influence both the intrinsic and externally induced spin polarizations in two-dimensional $d$-wave altermagnets. For the induced spin polarizations, we employ a Rashba continuum model with electron-phonon coupling (EPC) treated at the static-Holstein level and analyze the spin Edelstein effect using the Kubo linear-response formalism. We find that moderate-to-strong EPC progressively suppresses the induced polarization via both intraband and interband channels, with a critical coupling marking the onset of complete spin Edelstein depolarization. The depolarization transition arises from a phonon-induced energy renormalization that pushes the spin-split bands anisotropically above the chemical potential, leading to a complete collapse of the Fermi surface. While (de)polarization can occur even in the Rashba non-altermagnetic phase, it remains isotropic. The presence of altermagnetism makes it anisotropic and breaks the conventional antisymmetry between spin susceptibilities that occurs with pure spin-orbit coupling, rendering the effect highly relevant for spintronic applications. We further investigate how the phonon coupling to the altermagnetic order, Rashba spin-orbit strength, and carrier doping collectively tune the depolarization transition. Our findings demonstrate that phonon scattering (e.g., through various substrates) offers a powerful means for on-demand control of spin polarization, enabling reversible switching between spin-polarized and depolarized states—a key functionality for advancing spin logic architectures and optimizing next-generation spintronic devices.
I Introduction
Altermagnets are a recently identified class of magnetic materials that combine compensated magnetic order with spin-split electronic bands, despite having no net magnetization [1, 2, 3, 4, 5, 6, 7, 8]. Their unusual properties arise from symmetry-protected spin polarization and a mechanism of time-reversal symmetry breaking that preserves global spatial inversion. This, in turn, leads to momentum-dependent spin splittings and nontrivial spin textures even in the absence of external magnetic fields [9, 10, 11, 12]. The combination of vanishing macroscopic magnetization and strongly spin-polarized states makes altermagnets attractive for spintronic applications, where minimizing stray fields and energy dissipation is crucial [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. Among them, $d$-wave altermagnets have been mainly studied and noted to host unconventional current-driven spin phenomena [27, 14, 15, 28, 29, 30].
The capability to control intrinsic spin polarization and to externally induce spin polarization in altermagnets offers a powerful pathway for device integration in spintronics. One promising route is through the Edelstein effect [31, 32, 33, 34, 35, 36, 37, 38, 38, 35, 39, 40], where electric fields drive spin polarizations in systems with broken inversion symmetry. Such inversion asymmetry, in turn, can be engineered externally, for example by applying gate potentials that induce Rashba spin-orbit coupling (RSOC) [41, 42]. In this way, altermagnets can be electrically tuned to host spin-polarized states on demand, opening pathways toward low-power spin-charge interconversion, gate-controllable spin currents, and functional components for next-generation spintronic devices. Monolayer altermagnets have been shown to exhibit spin polarization when subjected to a perpendicular electric field [43]. In parallel, theoretical proposals have uncovered a rich variety of current-driven spin responses: relativistic mechanisms at altermagnetic interfaces [44], nonrelativistic ones across bulk altermagnets [39], $p$-wave anti-altermagnets [45, 46, 33], and magnets with chiral spin textures [40].
Although altermagnets offer exciting opportunities for spintronic applications, the impact of lattice dynamics (phonons) on their intrinsic and extrinsic spin polarizations is still largely unexplored. Electron-phonon coupling (EPC) can induce quasiparticle renormalization and enable phonon-assisted transport processes in both electronic and magnetic systems [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. Very recently, Irosh [58] demonstrated that electron pairing mediated by dispersive phonons in altermagnets can give rise to reentrant superconductivity as a function of temperature. Leraand et al. [59] further showed that, within the weak EPC regime, the dominant superconducting instability in altermagnets is odd in momentum and even in spin, hosting spin-polarized Cooper pairs. In addition, Steward et al. [60] reported that coupling the altermagnetic order to lattice vibrations leads to the emergence of a hybridized paramagnon-polaron mode. In another study, Hodt et al. [61] found that electron-phonon scattering processes can strongly enhance the spin conductivity of altermagnets at finite frequencies. He et al. [62] have also emphasized the pivotal role of electron-phonon scattering in shaping both the efficiency and the directionality of charge-to-spin conversion in altermagnets. Nonetheless, despite phonon’s ubiquity in electronic systems, the role of slow phonons in shaping the low-energy spin excitations and spin Edelstein effect in Rashba altermagnets has so far received little attention and remains entirely uncharted.
In typical two-dimensional (2D) Rashba systems, the electronic bandwidth far exceeds the characteristic phonon energy scale [63, 64], implying that phonons evolve on much slower timescales than electronic degrees of freedom. This separation of timescales forms the basis of many theoretical approaches to spin-related phenomena in altermagnets, which remain largely unexplored. Building on the above recent developments in understanding EPC effects in altermagnets, this work addresses the open question of how slow phonons, in conjunction with intrinsic spin polarization, affect the induced spin Edelstein polarizations in $d$-wave altermagnets. To capture this interplay, we construct a minimal continuum model at low temperatures that includes both RSOC and linear EPC (between lattice displacements and both spin-degenerate and spin-split states) at the static-Holstein level. Using the Kubo formalism, we compute induced spin Edelstein polarization over an extensive range of parameters, revealing distinctive fingerprints of the combined effects of lattice dynamics and the underlying crystalline symmetry in altermagnets.
Our analysis uncovers several key insights: (i) EPC induces a net magnetization, or equivalently an out-of-plane spin polarization, and renormalizes the intrinsic spin splitting in $d$-wave altermagnets; (ii) intraband and interband scattering processes strongly suppress the Edelstein response with increasing EPC, eventually leading to a critical depolarization point where the induced spin polarization vanishes permanently; (iii) the presence of altermagnetism enhances anisotropy and modifies the usual antisymmetry between spin susceptibilities, producing directionally dependent (de)polarization; and (iv) Rashba strength and doping provide tunable knobs that shift the onset of depolarization and govern whether spin polarization is stabilized.
Our findings establish phonon engineering as a promising route to optimize next-generation spintronic devices. Enhancing phonon scattering (e.g., using various substrates which produce different EPCs) can suppress nonequilibrium spin accumulation, reducing spin-related dissipation in spin-orbit torque devices. Moreover, on-demand switching between spin-polarized and depolarized states—crucial for spin logic architectures where spin rather than charge encodes information [65]—relies on controlled depolarization to reset logic elements, erase stored information, and prevent spin leakage.
The remainder of the paper is structured as follows. Section II introduces the model Hamiltonian, including the treatment of RSOC and EPC. In Sec. III, we outline the calculation of spin Edelstein polarizations. Section IV presents the results. Brief notes on the experimental feasibility of our results are given in Sec. V. Finally, Sec. VI summarizes key findings.
Figure 1: (a) Real-space representation of the staggered spin configuration on a square lattice of a pristine $d_{x^{2}-y^{2}}$-wave altermagnet with phonons included and without gating, where the spin-dependent orbital texture alternates between sublattices. Sublattice states are linearly coupled to slow Holstein phonons, depicted as shaded (vibrating) sites. (b) Corresponding band structure along high-symmetry paths near the $\Gamma$-point, displaying spin-split dispersions for spin-up and spin-down channels.
II Hamiltonian model
II.1 Pristine $d_{x^{2}-y^{2}}$-wave altermagnet
The low-energy quasiparticle excitations of a 2D Rashba-coupled $d_{x^{2}-y^{2}}$-wave altermagnet, Fig. 1(a), can be described by an effective two-band Hamiltonian acting in spin space. It consists of three main contributions: (i) a kinetic energy term, proportional to the identity matrix $\sigma_{0}$, which is spin-independent; (ii) an anisotropic spin-splitting term, proportional to $\sigma_{z}$, that originates from the $d$-wave altermagnetic order; and (iii) a RSOC term, proportional to $\sigma_{x}$ and $\sigma_{y}$, which breaks inversion symmetry, stemming from the gate electrodes [41, 42]. Explicitly, the pristine Rashba Hamiltonian reads
$\mathcal{H}_{\vec{k}}=\alpha_{\vec{k}}\,\sigma_{0}+\beta_{\vec{k}}\,\sigma_{z}+\lambda_{\rm R}\left(k_{y}\sigma_{x}-k_{x}\sigma_{y}\right)\,$, (1)
where $\lambda_{\rm R}$ denotes the RSOC strength. The momentum-dependent coefficients are given by
$\alpha_{\vec{k}}=\frac{\hbar^{2}}{2m}\,(k_{x}^{2}+k_{y}^{2})\quad{\rm and}\quad\beta_{\vec{k}}=\frac{\hbar^{2}\beta}{2m}\,(k_{x}^{2}-k_{y}^{2})\,$, (2)
with $m$ the effective electron mass and $0<\beta<1$ parametrizing the $d_{x^{2}-y^{2}}$-wave anisotropy of the spin splitting. In this model, $\beta_{\vec{k}}$ even in $\vec{k}$ ($\pi/a<k_{x}<\pi/a$ and $\pi/a<k_{y}<\pi/a$ with lattice constant $a$) preserves $C_{4}$ rotational symmetry, while RSOC term breaks inversion symmetry, thus, spin locks to momentum. Accordingly, RSOC wants to twist spin around the Fermi surface, while altermagnetism distorts it anisotropically. We briefly note that although our present study focuses on the $d_{x^{2}-y^{2}}$-wave anisotropy of the spin splitting, other symmetry patterns can be readily incorporated by rotating the plane of the altermagnet. For instance, a $45^{\circ}$ rotation of the plane yields a $d_{xy}$-wave altermagnet characterized by $\beta_{\vec{k}}=\frac{\hbar^{2}\beta}{m}\,k_{x}k_{y}$.
Diagonalization of Eq. (1) yields the band dispersions
$E_{\vec{k},s}=\alpha_{\vec{k}}+s\sqrt{\beta_{\vec{k}}^{2}+\lambda_{\rm R}^{2}k^{2}}\equiv\alpha_{\vec{k}}+s\,d_{\vec{k}},$ (3)
where $s=\pm$ labels the spin-resolved bands, $k=\sqrt{k_{x}^{2}+k_{y}^{2}}$, and $d_{\vec{k}}$ defines the magnitude of the effective pseudospin field. These two branches correspond to states in which the quasiparticle spin is either aligned (++) or anti-aligned (-) with the pseudospin texture generated by the combined $d$-wave anisotropy and Rashba coupling. The pristine low-energy band dispersion of $d_{x^{2}-y^{2}}$-wave altermagnet with $\lambda_{\rm R}=0$ is presented in Fig. 1(b). The momentum-dependent spin splitting is clearly visible. The eigenstates of Eq. (1) take the form of normalized spinors
$|\psi_{\vec{k},+}\rangle$ $=\frac{1}{\sqrt{2d_{\vec{k}}(d_{\vec{k}}+\beta_{\vec{k}})}}\begin{pmatrix}\beta_{\vec{k}}+d_{\vec{k}}\\ \\ \lambda_{\rm R}(k_{y}-ik_{x})\end{pmatrix}$, (4a)
$|\psi_{\vec{k},-}\rangle$ $=\frac{1}{\sqrt{2d_{\vec{k}}(d_{\vec{k}}+\beta_{\vec{k}})}}\begin{pmatrix}\lambda_{\rm R}(k_{y}+ik_{x})\\ \\ -\beta_{\vec{k}}-d_{\vec{k}}\end{pmatrix}$. (4b)
II.2 Phonon-dressed $d_{x^{2}-y^{2}}$-wave altermagnet
We now introduce a local, spinful Holstein-type electron-phonon interaction [66, 67], see Fig. 1(a), described by the phonon Hamiltonian
$\mathcal{H}_{\text{p}}=\sum_{i}\left(\frac{P_{i}^{2}}{2M}+\frac{1}{2}M\omega_{\rm p}^{2}Q_{i}^{2}\right)\,$, (5)
where $Q_{i}$ represents the displacement of the ion at site $i$, $P_{i}$ is the conjugate momentum, $M$ is the ion mass, and $\omega_{\rm p}$ is the characteristic phonon frequency.
In altermagnets at half filling, the spin-sublattice symmetry ensures that the electronic density is spatially uniform, thus, any static lattice distortion would break this symmetry without lowering the energy. To maintain this symmetry, we impose that the net force on the lattice vanishes at half filling. Concretely, this means that the phonon displacement $Q_{i}$ couples to density fluctuations rather than the absolute electron density, so that $\langle n_{i,\uparrow}+n_{i,\downarrow}\rangle=1$. This centers the lattice equilibrium configuration at half filling and ensures that the electron-phonon interaction vanishes in the absence of charge or spin-density fluctuations. This is a standard assumption in electron-phonon models [68, 69, 70, 71, 72, 73, 74] and corresponds to the standard local linear coupling. However, the intrinsic altermagnetic order of the sublattices allows us to in general incorporate spin-dependent coupling constants, so that the electron-phonon interaction is described by
$\mathcal{H}_{\text{e-p}}=\sum_{i}\left[g_{\uparrow}\left(n_{i,\uparrow}-\tfrac{1}{2}\right)+g_{\downarrow}\left(n_{i,\downarrow}-\tfrac{1}{2}\right)\right]Q_{i}\,.$ (6)
We stress that in our model, we subtract the half filling density $1/2$ rather than the instantaneous average density. In standard electron-phonon models with a quadratic phonon potential, this choice would only shift the equilibrium lattice configuration and could be absorbed into the chemical potential without changing the physics. However, in our case, the fixed chemical potential renders the choice physically meaningful: it guarantees that the coupling term vanishes at half filling where the lattice symmetry remains unbroken.
Crucially, since lattice displacements $Q_{i}$ are even under time reversal while $n_{i,\uparrow}-n_{i,\downarrow}$ is odd, the bilinear $(n_{i,\uparrow}-n_{i,\downarrow})Q_{i}$ is time-reversal-odd and therefore can appear only in a system that already breaks time-reversal symmetry. Although the displacement vectors vary in direction across different sites and spin states, their overall pattern follows a staggered arrangement: spin-up along $x$ experiences $Q_{i}$ while spin-down experiences $Q_{i}$, with the roles reversed along $y$. Accordingly, a nonzero $g_{\uparrow/\downarrow}$ is symmetry-allowed in altermagnets, but it is forbidden in time-reversal-symmetric crystals unless an external time-reversal-odd field is present. In fact, lattice distortions alter bond lengths and angles, and in altermagnets with intrinsically anisotropic spin textures, such geometric modulations generically induce spin-asymmetric corrections to the electronic structure, which at low energies project onto a local spin-dependent potential.
We therefore introduce symmetric and antisymmetric combinations of the coupling constants as
$g=\frac{g_{\uparrow}+g_{\downarrow}}{2},\quad\tilde{g}=\frac{g_{\uparrow}-g_{\downarrow}}{2}\,.$ (7)
We define the spin-dependent couplings as $g_{\uparrow}=g+\tilde{g}$ and $g_{\downarrow}=g-\tilde{g}$. Substituting these into the general EPC term in Eq. (6), we obtain
$\mathcal{H}_{\text{e-p}}=\sum_{i}\left[g\bigl(n_{i,\uparrow}+n_{i,\downarrow}-1\bigr)+\tilde{g}\bigl(n_{i,\uparrow}-n_{i,\downarrow}\bigr)\right]Q_{i}\,.$ (8)
This expression naturally separates into two contributions: a symmetric part coupling to the total local electron density $n_{i}=n_{i,\uparrow}+n_{i,\downarrow}$ and an antisymmetric part coupling to the local spin imbalance $m_{i}=n_{i,\uparrow}-n_{i,\downarrow}$. As previously noted, lattice displacements $Q_{i}$ modulate crystal fields and exchange/hopping in the presence of RSOC, producing a spin-asymmetric on-site potential. In a two-sublattice altermagnet the net magnetization induced by $\tilde{g}$ becomes finite for staggered (or otherwise symmetry-breaking) displacement patterns, yielding a macroscopic piezomagnetic response. Thus, microscopically, $\tilde{g}$ directly maps onto the piezomagnetic coefficient measurable in displaced altermagnets [75, 76, 77, 78, 60]. We emphasize that in the absence of altermagnetism, the effective coupling $\tilde{g}$ is forbidden, implying that the two parameters $\beta$ and $\tilde{g}$ are interdependent.
Assuming a staggered phonon displacement pattern across the sublattices but with a static-Holstein displacement amplitude, we focus on the regime of slow phonons where the lattice responds quasi-statically to the electronic environment, i.e., we set $\langle Q_{i}\rangle=Q_{0}$. The slow phonon-dressed Hamiltonian in momentum space can be written within the static-Holstein approximation as
$\mathcal{H}_{\vec{k}}^{\text{SH}}=\left(\alpha_{\vec{k}}+gQ_{0}\right)\sigma_{0}+\left(\beta_{\vec{k}}+\tilde{g}Q_{0}\right)\sigma_{z}+\lambda_{\rm R}\left(k_{y}\sigma_{x}-k_{x}\sigma_{y}\right).$ (9)
The phonon-induced terms $gQ_{0}$ and $\tilde{g}Q_{0}$ play distinct roles in the low-energy electronic structure. The shift $gQ_{0}$ renormalizes the chemical potential and, away from half filling, provides a route for phonon-assisted tuning of the carrier density. In contrast, the spin-dependent term $\tilde{g}Q_{0}$ acts as a dynamical Zeeman-like field, lifting the spin degeneracy and enabling transient spin polarization or spin-current control. When $\tilde{g}=0$, only the $gQ_{0}$ term is present, producing an overall band shift without modifying the spin texture arising from $\beta_{\vec{k}}$ or RSOC; the spectrum therefore remains isotropic along the $\Gamma\!\to\!X$ and $\Gamma\!\to\!Y$ directions. By contrast, a finite $\tilde{g}Q_{0}$ introduces opposite shifts for spin-up and spin-down bands, leading to spin-dependent anisotropies even though the coupling itself is momentum independent.
The equilibrium phonon displacement $Q_{0}$ is determined by minimizing the total free energy, taking into account both the harmonic restoring force of the lattice and the coupling to the electron density:
$M\omega_{\rm p}^{2}Q_{0}={}-g\,[n_{\rm e}-1]-\tilde{g}\,m\,$, (10)
where $n_{\rm e}=n_{\uparrow}+n_{\downarrow}$ is the total density and $m=n_{\uparrow}-n_{\downarrow}$ is the imbalance density, with
$n_{\rm e}={}$ $\frac{1}{N_{\vec{k}}}\sum_{\vec{k},s}f_{\vec{k},s}\langle\tilde{\psi}_{\vec{k},s}|\sigma_{0}|\tilde{\psi}_{\vec{k},s}\rangle\,$, (11a)
$m={}$ $\frac{1}{N_{\vec{k}}}\sum_{\vec{k},s}f_{\vec{k},s}\langle\tilde{\psi}_{\vec{k},s}|\sigma_{z}|\tilde{\psi}_{\vec{k},s}\rangle\,$, (11b)
where $\tilde{\psi}$ is phonon-dressed spinor, $N_{\vec{k}}$ denotes the number of $\vec{k}$-points in the Brillouin zone, and $f_{\vec{k},s}$ is the Fermi-Dirac distribution function, $f_{\vec{k},s}=(e^{(\mathcal{E}_{\vec{k},s}-\mu)/k_{\rm B}T}+1)^{-1}$; $\mu$ is the bare chemical potential and $\mathcal{E}_{\vec{k},s}$ is the phonon-dressed band dispersion:
$\mathcal{E}_{\vec{k},s}=\alpha_{\vec{k}}+gQ_{0}+s\sqrt{\left(\beta_{\vec{k}}+\tilde{g}Q_{0}\right)^{2}+\lambda_{\rm R}^{2}k^{2}}\equiv\tilde{\alpha}_{\vec{k}}+s\,\tilde{d}_{\vec{k}}.$ (12)
The phonon-dressed quasiparticle dispersion inherits several key features from the underlying electronic structure and the EPC. The first term, $\tilde{\alpha}_{\vec{k}}=\alpha_{\vec{k}}+gQ_{0}$, represents a uniform shift of all bands due to the static lattice distortion $Q_{0}$. The second term, $\tilde{d}_{\vec{k}}=\sqrt{(\beta_{\vec{k}}+\tilde{g}Q_{0})^{2}+\lambda_{\rm R}^{2}k^{2}}$, controls the spin splitting between $s=\pm 1$ bands. Importantly, $\beta_{\vec{k}}+\tilde{g}Q_{0}$ is an even but anisotropic function of momentum: the coupling $\tilde{g}Q_{0}$ modifies its magnitude in a direction-dependent way, preserving inversion symmetry but breaking $C_{4}$ rotational symmetry. As a result, the dressed bands show spin-dependent gaps and anisotropic distortions whose strength grows with $Q_{0}$, explaining the deformation. The Rashba term $\lambda_{\rm R}^{2}k^{2}$ further broadens the spin splitting, ensuring that even when $\beta_{\vec{k}}+\tilde{g}Q_{0}\to 0$ along some directions, the spin bands remain separated by a gap proportional to $|\lambda_{\rm R}k|$.
Since our effective model is valid only within a restricted energy window, we focus on very low temperatures, ensuring that thermal excitations remain within the model’s regime of validity. In the zero-temperature limit, the Fermi-Dirac occupations reduce to $f_{\vec{k},s}\approx\Theta(\mu-\mathcal{E}_{\vec{k},s})$, where $\Theta$ is the Heaviside function. Therefore, $n_{\rm e}$ and $m$ satisfy the self-consistent equations
$n_{\rm e}\approx{}$ $\frac{1}{N_{\vec{k}}}\sum_{\vec{k}}\left[\Theta\left(\mu-\mathcal{E}_{\vec{k},+}(Q_{0})\right)+\Theta\left(\mu-\mathcal{E}_{\vec{k},-}(Q_{0})\right)\right]\,$, (13a)
$m\approx{}$ $\frac{1}{N_{\vec{k}}}\sum_{\vec{k}}\left[\Theta\left(\mu-\mathcal{E}_{\vec{k},+}(Q_{0})\right)-\Theta\left(\mu-\mathcal{E}_{\vec{k},-}(Q_{0})\right)\right]\,$, (13b)
where $Q_{0}$ depends on both $n_{\rm e}$, and $m$.
III Spin Edelstein effect
In crystals lacking inversion symmetry, spin-orbit interaction allows an external electric field to generate a net nonequilibrium spin or orbital polarization. When it involves electron spin, this effect is referred to as the spin Edelstein effect or Rashba-Edelstein effect [31, 35, 36]. It is important to distinguish between “spin splitting” and “spin polarization” in the context of our study. Spin splitting refers to the intrinsic separation of energy bands for opposite spin projections, which arises from the altermagnetic order and/or RSOC. It is encoded directly in the band structure and exists even in the absence of external perturbations. In contrast, spin polarization denotes an induced magnetization, which arise from electron-phonon interactions in our study. Unlike spin splitting, spin polarization is not a fundamental property of the electronic structure but rather reflects the population imbalance between spin states, which can be strongly modulated by lattice distortions or other interactions. Within the linear response regime, the induced expectation value of a spin operator $\hat{S}^{\ell}$ ($\ell\in\{x,y,z\}$) due to an in-plane electric field $\mathbb{E}^{j}$ along direction $j\in\{x,y\}$ can be written as [37, 38]
$\langle\hat{S}^{\ell}\rangle=\sum_{j}\chi_{\ell j}\mathbb{E}^{j}\,$, (14)
where $\chi_{\ell j}$ defines the linear susceptibility, also called the Edelstein response tensor. The response tensor $\chi_{\ell j}$ can be computed using the Kubo formalism. It contains contributions from intraband processes within a single band as well as from interband coherences, given by
$\chi_{\ell j}={}$ $\frac{e}{4\pi^{2}}\int d^{2}k\Bigg[\tau_{\rm intra}\,\sum_{s}\frac{\partial f_{\vec{k},n}}{\partial E_{\vec{k},s}}S^{\ell}_{\vec{k},ss}\,v^{j}_{\vec{k},ss}$
$-i\sum_{s\neq s^{\prime}}\frac{f_{\vec{k},s}-f_{\vec{k},s^{\prime}}}{E_{\vec{k},s}-E_{\vec{k},s^{\prime}}}\frac{S^{\ell}_{\vec{k},s^{\prime}s}\,v^{j}_{\vec{k},ss^{\prime}}}{E_{\vec{k},s^{\prime}}-E_{\vec{k},s}+i/\tau_{\rm inter}}\Bigg]\,.$ (15)
In this study, we set the intraband and interband scattering rates to be equal, $\tau^{-1}_{\rm intra}=\tau^{-1}_{\rm inter}=0.5~\mathrm{eV}$, a value consistent with typical metallic systems [38, 35]. The operator matrix elements are defined as
$S^{\ell}_{\vec{k},s^{\prime}s}={}$ $\frac{\hbar}{2}\langle\tilde{\psi}_{\vec{k},s^{\prime}}|\hat{\sigma}^{\ell}|\tilde{\psi}_{\vec{k},s}\rangle\,$, (16a)
$v^{j}_{\vec{k},ss^{\prime}}={}$ $\frac{1}{\hbar}\langle\tilde{\psi}_{\vec{k},s}|\partial\mathcal{H}_{\vec{k}}^{\text{SH}}/\partial k_{j}|\tilde{\psi}_{\vec{k},s^{\prime}}\rangle\,.$ (16b)
With the phonon-dressed Hamiltonian from Eq. (9), the corresponding matrix elements can be computed. In the zero-temperature limit, the derivative of the Fermi function reduces to a Dirac delta $\frac{\partial f}{\partial E}\to-\delta(E-\mu)$. Substituting this into the intraband term yields a simplified expression for the low-temperature response $\chi^{\rm intra}_{\ell j}=-\tfrac{e}{4\pi^{2}}\tau_{\rm intra}\int d^{2}k\sum_{s}\delta(E_{\vec{k},s}-\mu)\,S^{\ell}_{\vec{k},ss}\,v^{j}_{\vec{k},ss}$. This form directly connects the intraband Edelstein response to the states at the Fermi surface and their corresponding operator and velocity matrix elements.
In the following, we compute the spin Edelstein effect within our phonon-dressed model, which accounts for the slow lattice dynamics. This allows us to investigate how the phonon-induced renormalization of the chemical potential, $gQ_{0}$, and the generation of an effective magnetization, $\tilde{g}Q_{0}$, influence the spin response. In particular, we analyze how these phonon-mediated modifications compete or cooperate with other intrinsic parameters of the model to shape the magnitude and orientation of the induced spin polarization. In our simulations, for slow phonon, we take the limits $M\to\infty$ and $\omega_{\rm p}\to 0$ such that the product $M\omega_{\rm p}^{2}$, corresponding to the stiffness of the phonon potential, is fixed to unity.
[…]
IV Results and discussion
We summarize the key parameters in our model as follows: $g$ (EPC to on-site electron densities), $\tilde{g}$ (EPC to staggered electron densities), $0<\beta<1$ (altermagnetic strength), $\lambda_{\rm R}$ (RSOC strength), and $\mu$ (chemical potential). For the EPCs, we apply the Lindemann criterion [79], which limits atomic displacements to at most 10% of the lattice constant to prevent lattice melting. Specifically, for a representative typical lattice constant of $4-5\,\text{\AA }$ in target $d$-wave altermagnets, we select $g$ such that the equilibrium displacement $Q_{0}$ does not exceed $0.5\,\text{\AA }$. Furthermore, we set $\tilde{g}<g$, since the staggered energy responsible for spin splitting is inherently smaller than the on-site hopping energy. Throughout the subsequent analyses, we set $\tilde{g}=g/4$ unless stated otherwise. Finally, $\lambda_{\rm R}$ and $\mu$ are tuned to cover the bandwidth of the spin channels, ensuring that the spin response is properly aligned with the electronic structure.
In the numerical simulations, we iteratively solve Eqs. (13a) and (13b) for the given EPCs and bare chemical potential $\mu$ [80]. Once $n_{\rm e}$ and $m$ are specified, the equilibrium phonon displacement is given by $Q_{0}=-(g\,[n_{\rm e}-1]+\tilde{g}\,m)/M\omega_{\rm p}^{2}$. Even though $Q_{0}$ appears as a static-Holstein model artifact, it reflects a stable minimum of the full Hamiltonian. It does not decay to zero or diverge, demonstrating that the finite displacement is self-consistent and dynamically robust. In the limit of extremely soft phonon modes ($\omega_{\rm p}\to 0$), the phonon dynamics become much slower than the electronic timescales. Consequently, the kinetic term $P_{i}^{2}/2M$ vanishes, rendering the phonon field effectively static. Our approach can be viewed as a self-consistent, quasi-static treatment of the electron-phonon coupled system, in which the phonon displacement $Q_{0}$ is determined by minimizing the full Hamiltonian for given electronic densities. Unlike a full Holstein model solution, we do not average over multiple phonon configurations or include dynamical phonon fluctuations, and thus many-body effects such as the frequency-dependent self-energy are neglected. While this corresponds formally to the zero-frequency limit of the Migdal perturbation theory at lowest order [81], it is more precise to describe our method as a self-consistent static-Holstein approximation, capturing non-perturbative, density-driven lattice shifts beyond standard linearized treatments.
Once the complete Hamiltonian in Eq. (9) is established, we can apply the Kubo formula from Eq. (III) to calculate the spin Edelstein effect. This approach uses the phonon-dressed Hamiltonian’s eigenvalues and eigenstates to determine the linear response of the spin polarization to an applied electric field.
Figure 2(a) illustrates the evolution of the spin Edelstein susceptibilities, $\chi_{\ell j}/\chi_{0}$, where $\chi_{0}=e/4\pi^{2}$ sets the reference scale, with increasing EPC strength $g$ at $\mu=0$. For weak EPC ($g\lesssim 0.06$ eV/Å), the susceptibilities remains nearly constant. In this regime, electron-phonon scattering is too weak to significantly modify the spin-split band structure, which reflects the perturbative regime. In other words, the spin-momentum locking characteristic of Rashba altermagnet ensures a robust spin Edelstein response in this regime. To explicitly demonstrate the breaking of the antisymmetry relation in the spin Edelstein response, $\chi_{\ell j}\neq-\chi_{j\ell}$, we plot the quantity $\chi_{j\ell}$ for comparison. If the antisymmetric condition $\chi_{\ell j}=-\chi_{j\ell}$ were preserved under the influence of EPC, the curves corresponding to $\chi_{\ell j}$ and $\chi_{j\ell}$ would coincide, exhibiting no deviation. We note that, owing to the symmetry of our model—specifically, $v^{z}_{{\vec{k}},ss^{\prime}}=0$ arising from the absence of a $k_{z}$ component in the Hamiltonian—the spin Edelstein effect cannot generate out-of-plane ($z$-) polarization, leading to $\chi_{\ell z}=\chi_{zj}=0$. Thus, only in-plane spin polarizations are induced.
For all $g$ values, the dominant contributions arise from the off-diagonal elements $\chi_{xy}$ and $\chi_{yx}$. These terms encode the transverse Edelstein response, reflecting the strong spin-momentum locking induced by RSOC. In contrast, the diagonal components $\chi_{xx}$ and $\chi_{yy}$ remain much smaller, with $\chi_{xx}$ slowly decreasing and $\chi_{yy}$ staying nearly constant across a large range of $g$. As $g$ increases, both $\chi_{xy}$ and $\chi_{yx}$ undergo a monotonic decrease, eventually touching zero near $g_{\rm c}\approx 0.25$ eV/Å. This zero response signals the point of spin Edelstein depolarization.
The distinct vanishing of $\chi_{xy}$ and $\chi_{yx}$ indicates that the depolarization affects both directions of the transverse spin response anisotropically and can be rationalized from symmetry considerations in the phonon-renormalized band structure. From Eq. (12), the anisotropic term $\beta_{\vec{k}}+\tilde{g}Q_{0}$ modifies the spin-split eigenvalues differently along the $\Gamma\!\to\!X$ and $\Gamma\!\to\!Y$ directions. Since $\chi_{xy}$ and $\chi_{yx}$ are sensitive to the momentum-resolved band dispersions, the anisotropic shift induced by the phonon term causes the conditions for depolarization (i.e., when certain Fermi states no longer contribute) to occur at different EPC strengths along these directions. In other words, the symmetry-breaking effect of $\tilde{g}Q_{0}$ on $\tilde{d}_{\vec{k}}$ lifts the degeneracy between $x$ and $y$ susceptibilities, naturally leading to distinct vanishing points for $\chi_{xy}$ and $\chi_{yx}$ (see Fig. 4 for confirmation).
Figure 2(b) provides further insight by decomposing the transverse component $\chi_{xy}/\chi_{0}$ into its intraband and interband contributions. A similar analysis can be carried out for the other components. The intraband part dominates at all coupling strengths, reflecting the usual Edelstein mechanism in which current-carrying states at the Fermi surface induce spin polarization. The interband part remains in relative weight with $g$.
To understand the depolarization phenomenon, we examine how the spin-resolved band structure evolves in the presence of slow static-Holstein phonon. Specifically, we revisit the expression for the spin susceptibility in Eq. (III). The vanishing of $\chi_{\ell j}$ requires the velocity term $v^{j}_{\vec{k}}$ to become zero over the dominant regions of the Brillouin zone. A straightforward scenario arises when the spin-split bands are shifted away from the zero Fermi energy due to $\mu=0$ in our simulations. In such cases, the group velocity $v^{j}_{\vec{k}}=\hbar^{-1}\partial\mathcal{E}_{\vec{k}}/\partial k_{j}$ approaches zero due to the absence of bands near the Fermi energy, rendering the electrons effectively immobile in response to an applied electric field. Consequently, the charge current—and hence the current-induced spin polarization—vanishes.
To identify the critical coupling strength $g_{\rm c}$ at which the spin Edelstein polarization is fully suppressed, marking the onset of depolarization, we consider the band dispersion $\mathcal{E}_{\vec{k},s}$. Since the lower band (s=1s=-1) gives the smallest energies, the condition for the absence of zero-energy states is $\min_{\vec{k}}[\mathcal{E}_{\vec{k},-}]>0$.
Figure 3 shows the evolution of the minimum of the lower spin-split band as the EPC $g$ is increased. The critical coupling $\tilde{g}_{\rm c}$ emerges when $\min_{\vec{k}}[\mathcal{E}_{\vec{k},-}]=0$, with its value depending on the specific parameter space. In Fig. 3(a), increasing the altermagnetic anisotropy $\beta$ shifts the curves downward, indicating that stronger anisotropy delays the depolarization transition to higher $\tilde{g}_{\rm c}$. Similarly, in Fig. 3(b), increasing the RSOC $\lambda_{\rm R}$ lowers the minimum band energy, thereby increasing $\tilde{g}_{\rm c}$ and delaying the onset of depolarization. The circles mark the critical coupling $\tilde{g}_{\rm c}$ for each parameter set, defined by the condition $\min[\mathcal{E}_{\vec{k},-}]=\mu$. A slight mismatch arises between $\tilde{g}_{\rm c}$ and the actual onset of depolarization at $g_{\rm c}$ in the Edelstein response. This discrepancy originates from the fact that depolarization sets in only after the zeros of $\min[\mathcal{E}_{\vec{k},-}]$ are crossed, i.e., when the Fermi surface is fully collapsed due to the EPC-induced energy shifts and band anisotropies. However, the EPC contribution determined by the condition $\min[\mathcal{E}_{\vec{k},-}]=\mu$ corresponds to the point where the minimal band energy just reaches the chemical potential, thus, $g_{\rm c}\gtrsim\tilde{g}_{\rm c}$.
Figure 3: Minimum of the lower spin-split band, $\min[\mathcal{E}_{\vec{k},-}]$, as a function of the EPC $g$ for two parameter variations. (a) Variation of the altermagnetic anisotropy $\beta$ and (b) variation of the RSOC strength $\lambda_{\rm R}$. The horizontal dashed line marks the zero Fermi energy $\mu$. Open circles indicate the critical EPC $\tilde{g}_{\rm c}$ at which the lower band touches $\mu$, marking the onset of spin Edelstein depolarization at $g_{\rm c}\gtrsim\tilde{g}_{\rm c}$.
While the absence of bands or the (an)isotropy of bands near the Fermi energy can be directly inferred from the electronic band structure, additional insights into the effects of the parameters can be obtained by examining the electronic density of states (DOS) in the presence of EPC. In particular, the separation of van Hove singularities along different directions emerges once the spin-resolved band structure becomes anisotropic due to variations in the model parameters. The DOS per unit area is defined as $\mathcal{D(E)}=\frac{1}{(2\pi)^{2}}\sum_{s=\pm}\int d^{2}k\,\delta\!\bigl(\mathcal{E}-\mathcal{E}_{{\vec{k}},s}\bigr)$. In polar coordinates $d^{2}k=k\,dk\,d\theta$, this becomes
$\mathcal{D(E)}={}\frac{1}{(2\pi)^{2}}\sum_{s=\pm}\int_{0}^{2\pi}d\theta\int_{0}^{\infty}k\,dk\delta\!\Bigl(\mathcal{E}-\tilde{\alpha}_{k,\theta}-s\,\tilde{d}_{k,\theta}\Bigr).$ (17)
To evaluate the $k$-integral, for each band $s$ we solve the root equation $\mathcal{E}=\tilde{\alpha}_{k,\theta}+s\,\tilde{d}_{k,\theta}$ for positive solutions $k_{i}$. For a simple isolated root $k_{i}$ the delta function gives a Jacobian factor, so the contribution of that root is $k_{i}/(2\pi|\frac{d}{dk}\bigl(\tilde{\alpha}_{k,\theta}+s\,\tilde{d}_{k,\theta}\bigr)\bigl|_{k=k_{i}}|)$. Therefore, the practical evaluation formula becomes
$\mathcal{D(E)}=\frac{1}{2\pi}\sum_{s=\pm}\sum_{k_{i}>0}\frac{k_{i}}{\bigg|\tilde{\alpha}^{\prime}_{k_{i}}+s\,\frac{\lambda_{\rm R}^{2}k_{i}}{\tilde{d}_{k_{i}}}+s\,\frac{\bigl(\beta_{k_{i}}+\tilde{g}Q_{0}\bigr)\,\partial_{k}\beta_{k}\bigl|_{k=k_{i}}}{\tilde{d}_{k_{i}}}\bigg|}\,$, (18)
where $\tilde{\alpha}^{\prime}_{k_{i}}=\partial_{k}\tilde{\alpha}_{k_{i}}$. Our static-Holstein approach neglects the dynamical electron-phonon self-energy, which in conventional systems with finite $\omega_{\rm p}$ produces kinks in the DOS at $|\mathcal{E}|=\omega_{\rm p}$. In the $\omega_{\rm p}\to 0$ limit, these effects approach the Fermi surface but are absent here, as they stem from the dynamical, not static, self-energy. It should be noted that in the above summation, in order to highlight the anisotropic modifications of the band structure, we treat $k_{i}$ separately along the $\Gamma\to X$ and $\Gamma\to Y$ paths.
[…]
Figure 4 presents the DOS $\mathcal{D}(\mathcal{E})$ along the high-symmetry $\Gamma\to X$ and $\Gamma\to Y$ directions for various values of the EPC $g$. At weak coupling [Fig. 4(a), $g=0.04$ eV/Å], the band dispersions along $X$ and $Y$ are highly symmetric, leading to overlapping isotropic DOS peaks. The sharp peaks in $\mathcal{D}(\mathcal{E})$ correspond to van Hove singularities originating from nearly degenrate states in the band structure. This, in turn, implies that the spin Edelstein susceptibilities satisfy $\chi_{\ell j}=-\chi_{j\ell}$, as illustrated by the spin Edelstein responses in Fig. 2(a) near $g=0.04$ eV/Å, reflecting the intrinsic antisymmetric character of the bands. Upon increasing $g$ to $0.12$ eV/Å  [Fig. 4(b)], a small anisotropy emerges between the $X$ and $Y$ directions, offering slightly antisymmetric breaking feature $\chi_{\ell j}\neq-\chi_{j\ell}$, as shown by the responses in Fig. 2(a) near $g=0.12$ eV/Å.
At $g\approx 0.18$ eV/Å  [Fig. 4(c)], the band dispersions become markedly asymmetric associated with more split peaks. The inset shows that the bands become slightly warped: the spin-down dispersion along $X$ shifts downward relative to $Y$. This anisotropy is directly reflected in the DOS, where a double-peak structure begins to form, indicating the splitting of van Hove singularities associated with anisotropic band edges. The critical EPC, above which permanent depolarization occurs, is $g_{\rm c}\approx 0.25$ eV/Å. For $g=0.28$ eV/Å  [Fig. 4(d)], the DOS peaks are more separated than in the previous case shown in Fig. 4(c). Due to the absence of spin bands at the Fermi energy (see the inset showing the spin-resolved band structure in Fig. 4(d)), the intraband and interband contributions to the spin Edelstein susceptibility in Eq. (III) are strongly suppressed. Thus, electrons are highly localized; they do not propagate easily, which can strongly suppress transport-related properties.
Thereby, the Edelstein depolarization occurs as the phonon-induced energy shift $gQ_{0}$ elevates the electronic bands above the chemical potential, effectively collapsing the Fermi surface. The detailed dependence of this depolarization on system parameters reflects the intrinsic nonlinearity of the static-Holstein treatment that determines the equilibrium displacement $Q_{0}$.
A key point worth asking is whether phonon-induced suppression or complete vanishing of the spin Edelstein effect can be effectively harnessed to control spin transport in practical spintronic applications. Our analysis indicates that controlled phonon scattering offers precisely this opportunity. When phonon interactions are moderate, they act to dampen excess spin polarization, thereby minimizing spin-related dissipation in spin-orbit torque devices. In contrast, under strong phonon-mediated spin depolarization is central to the operation of spin logic circuits [65]. Spin-polarized states represent the binary logic levels required for data encoding and signal transmission, while spin depolarization plays an equally critical role by resetting logic units, erasing residual spin memory, and electrically isolating different parts of the circuit to prevent spin leakage.
Since the (de)polarization occurs primarily in the dominant transverse susceptibilities, and given that $\chi_{\ell j}\neq-\chi_{j\ell}$ but behaves similarly to $\chi_{j\ell}$, we henceforth focus on $\chi_{xy}$ and analyze its evolution across the parameter space in the presence of EPC.
IV.1 Altermagnetic order effect
Next, we analyze how altermagnetism affects the persistence of spin Edelstein depolarization, allowing us to contrast the behavior with and without an intrinsic spin-split band structure. Figure 5(a) shows the dependence of the transverse spin Edelstein susceptibility $\chi_{xy}/\chi_{0}$ on the EPC strength $g$ for different values of the staggered lattice potential parameter, or equivalently, the altermagnetic order $\beta$. For $\beta=0$, the system also exhibits a gradual suppression of $\chi_{xy}/\chi_{0}$ as $g$ increases because it is well established that RSOC alone generates a finite spin Edelstein polarization as a direct consequence of inversion-symmetry breaking. Importantly, the depolarization transition, signaled by the zero of $\chi_{xy}/\chi_{0}$, still occurs even in the absence of altermagnetism.
[…]
Although the model inherently predicts an isotropic spin Edelstein polarization at $\beta=0$, corresponding to antisymmetric spin susceptibilities, we explicitly validate this behavior by examining the electronic band structure and DOS as in the preceding analyses. Figure 5(b) offers further insight by displaying the electronic DOS for $\beta=0$ and $g=0.2$ eV/Å, along with the corresponding electronic band structure $\mathcal{E}_{\vec{k}}$ along the high-symmetry directions $Y\leftarrow\Gamma\to X$ (inset). As illustrated, the DOS along both the $X$ and $Y$ directions exhibits identical features at $\beta=0$, confirming that in the absence of altermagnetic order, the induced spin Edelstein (de)polarization remains highly isotropic. By contrast, the introduction of a finite altermagnetic order parameter $\beta$, see Fig. 2, breaks this isotropy by inducing anisotropic spin-splitting in the band structure, thereby rendering the spin Edelstein response directionally dependent. This anisotropy is a crucial attribute for spintronic applications, where control over the spin polarization along specific crystallographic directions is often desired.
As $\beta$ increases, the induced spin polarization decreases and the depolarization point shifts slightly toward larger critical EPC strengths $g_{\rm c}$, indicating that the altermagnetic order delays the onset of depolarization. This was already confirmed by Fig. 3(a). Physically, this can be attributed to the enhanced band structure induced by the combined effects of EPC and altermagnetic order, which strongly suppresses the intraband and interband contributions to the susceptibility.
It should be noted that in conventional spin-orbit coupled systems, the Edelstein effect produces a nonequilibrium spin polarization that is subsequently reduced by relaxation processes, e.g. Dyakonov–Perel-type dephasing from momentum scattering in $\vec{k}$-dependent spin-orbit fields [82, 83] or Elliott–Yafet scattering from impurity and phonon interactions [84, 85]. These mechanisms dynamically limit the lifetime of spin accumulation but do not break the underlying asymmetry in induced polarizations. By contrast, the Edelstein (de)polarization in altermagnets breaks this asymmetry: opposite spin textures on different parts of their Fermi surface change upon Brillouin-zone integration, yielding anisotropic (de)polarizations even in the absence of scattering. This makes the effect intrinsic and symmetry-driven, rather than extrinsic and relaxation-driven.
IV.2 Phonon-altermagnetic order coupling effect
As the next investigation, we examine how EPC to the altermagnetic order, $\tilde{g}/g$, affects the persistence of spin Edelstein depolarization. It is observed, in Fig. 6(a), that for $\tilde{g}/g<1$, the rate of polarization suppression depends only weakly on $\tilde{g}/g$. Larger values of the ratio (but still below 1) lead to a slower increase in $\chi_{xy}/\chi_{0}$ and slightly shift the zero to larger $g$, implying that a stronger coupling to the staggered potential slightly enhances the susceptibility renormalization. At $\tilde{g}/g=1$, there is a strong shift of the spin bands, such that a much larger EPC is required to suppress the spin Edelstein polarization, as the intraband and interband transitions still contribute significantly to the susceptibility. This, in turn, means that stronger $\tilde{g}/g$ stabilizes the transverse polarization more efficiently.
[…]
We further highlight that in the absence of explicit EPC to the staggered potential, depolarization still occurs. However, in Fig. 6(b) at $g=0.2$ eV/Å, our DOS analysis and band structure (inset of Fig. 6(b)) reveal that the induced polarization and depolarization remain isotropic at $\tilde{g}/g=0$, since the DOS is identical along the $X$ and $Y$ directions of the Brillouin zone. The parameters $\tilde{g}$ and $\beta$ are intrinsically linked, such that changes in one necessarily influence the other, reflecting their mutual dependence within the system. Similar to the case of $\beta=0$, such isotropy is undesirable for spintronic applications, where directional control of spin responses is essential. In contrast, once the ratio $\tilde{g}/g$ is introduced, the polarization and depolarization become anisotropic (see Fig. 2), providing the anisotropic spin responses that are precisely what spintronics aims to exploit.
[…]
IV.3 Rashba spin-orbit coupling effect
[…]
Physically, an increase in the RSOC enhances both the spin-momentum locking and the effective spin splitting of the bands, which in turn tends to preserve finite spin states at the Fermi surface. This makes the current-induced spin polarization more robust against EPC-induced band renormalization, thereby delaying the suppression of the current-induced spin polarization.
[…]
IV.4 Electron and hole doping effects
Finally, we examine the case of the doping effect, in which the chemical potential is shifted away from the Fermi energy ($\mathcal{E}=0$ eV). Figure 8 demonstrates the influence of the chemical potential $\mu$ on the transverse spin Edelstein susceptibility $\chi_{xy}/\chi_{0}$ as the EPC strength $g$ is varied.
When the chemical potential is shifted below the band edge ($\mu=-0.02$ and $0.04~\mathrm{eV}$) due to hole doping, Fig. 8(a), the magnitude of $\chi_{xy}$ is reduced compared to the $\mu=0$ case, indicating that fewer carriers at the Fermi level participate in the Edelstein response. Interestingly, as $\mu$ decreases, depolarization occurs at smaller EPC strengths. This again originates from the reshaping of the Fermi surface in the presence of both Rashba coupling and electron-phonon renormalization.
For positive chemical potentials ($\mu=+0.02$ and $0.04~\mathrm{eV}$), corresponding to electron doping, the behavior shown in Fig. 8(b) differs significantly. The susceptibility curve exhibits a nonmonotonic dependence on $g$, characterized by an intermediate enhancement before eventually decaying. The peak in the susceptibility can be understood in the same way as in Fig. 7(b): for $\mu>0$ and $g>0.15$ eV/Å, an increasing number of states participate in intraband and interband transitions, leading to a temporary increase in susceptibility. Beyond this point, as the number of available states diminishes, the susceptibility begins to decrease. Thus, a strong particle-hole asymmetry appears in the Edelstein response under finite doping.
V Experimental perspective
Before concluding, we comment on the experimental feasibility of our predictions. A promising route to experimentally probe the proposed effects is to interface a 2D $d$-wave altermagnet—such as thin films of KV2Se2[86], RbV2Te2[87], RuO2 (noting that the altermagnetic character of RuO2 remains under active investigation) [88, 89, 14, 90], and $\kappa$-Cl [91]—with an appropriate substrate at low temperatures to induce lattice vibrations in the altermagnet layer. Crucially, since the underlying phonon-mediated mechanism requires only a linear coupling between lattice displacements and spin states, the proposed effect should be realizable across a broad class of altermagnetic thin films. Phonons at nominally zero temperature can also be generated with inelastic neutron scattering providing a direct route: incident neutrons transfer both momentum and energy to the lattice, populating phonon states even at zero temperature [92, 93].
Electrostatic control over the RSOC can be achieved using top and bottom gate electrodes separated by high-quality dielectric layers. Complementary to phonon generation and gating, time- and angle-resolved photoemission spectroscopy enables direct visualization of spin textures, band splittings, and symmetry-breaking phenomena [94]. Additionally, the Edelstein effect can be probed in Hall-bar geometries through electrical transport measurements.
We note, however, that our analysis relies on a low-energy static-Holstein model with a simplified, non-dispersive phonon mode. While it captures the essential physics of EPC-driven spin (de)polarization, effects arising from full-band structures, dispersive phonons, or beyond static-Holstein interactions may quantitatively or qualitatively modify the results. Accordingly, the present findings should be regarded as minimal-model predictions.
VI Summary and outlook
In this work, we have examined how slow lattice vibrations shape spin responses in $d$-wave altermagnets, systems that combine vanishing net magnetization with spin-split electronic bands. By formulating the problem within a Rashba continuum model coupled to slow phonons through a Holstein-type interaction, and treating the EPC at the static-Holstein level, we have tracked the evolution of the Edelstein effect over a broad range of parameters. The Kubo formalism provided access to distinguish the specific contributions of intraband and interband processes.
A central outcome of our study is that increasing EPC gradually suppresses the Edelstein spin polarization. Once the coupling exceeds a critical threshold, the system undergoes a spin Edelstein depolarization. At this threshold, the spin polarization collapses to zero permanently, signaling a qualitative restructuring of the spin channels. This regime is characterized by the disappearance of both intraband and interband transitions at zero Fermi energy. The (an)isotropic character of the suppressed polarization and depolarization strongly depends on the presence of altermagnetism. Lacking altermagnetic order, the effect is isotropic, whereas the interplay with altermagnetism produces pronounced anisotropy and eliminates the usual antisymmetry between spin susceptibilities. This finding suggests that altermagnetism is a key ingredient in realizing directionally selective spin responses, thereby enhancing prospects for spintronic applications. Our parameter-space analysis further revealed that the onset of depolarization is tunable. The critical EPC shifts with the strength of coupling to a staggered lattice potential and the Rashba interaction. Doping also plays a decisive role: moving the chemical potential away from the zero Fermi energy modifies the threshold EPC for depolarization. Negative chemical potentials (hole doping) reduce the susceptibility and promote faster depolarization, whereas positive chemical potentials (electron doping) enhance the susceptibility and delay the onset of depolarization.
In practice, phonons can be controlled through approaches such as strain engineering or substrate selection. The selective excitation of specific vibrational modes modulates the EPC and local crystal fields, thereby influencing the induced spin polarization. Thus, our study shows that moderate phonon scattering can be utilized to regulate spin-related dissipation in spin-orbit torque devices by attenuating excess spin polarization. In contrast, when strong phonon-mediated depolarization dominates, it enables efficient clearing of residual spin information, reinitialization of logic elements for subsequent operations, and electrical isolation of different circuit sections to prevent spin leakage, thereby providing a practical route for on-demand switching between spin-polarized and spin-neutral states. Together, these mechanisms provide complementary strategies for enhancing the efficiency and design flexibility of emerging spintronic technologies.
Looking ahead, several intriguing directions emerge for future investigations. First, exploring the interplay of phonon-induced spin depolarization with other types of spin-orbit interactions, such as Dresselhaus or proximity-induced effects, could reveal richer anisotropic spin responses. Second, extending these studies to multilayer altermagnets or heterostructures may enable enhanced tunability through interlayer coupling and interface engineering. Third, time-resolved experiments, including ultrafast optical or terahertz pump-probe techniques, could probe the dynamical evolution of the Edelstein effect under controlled phonon excitations, providing direct verification of our predictions. Fourth, combining phonon engineering with strain, gating, or substrate manipulation could allow precise control over spin polarization, opening pathways for optimized spintronic device design. Finally, extending these concepts to other material classes, such as topological insulators, 2D magnets, or Rashba systems with strong electron correlations, may reveal universal features of phonon-controlled spin responses. We leave these topics to be explored in future work.
Acknowledgments
M. Y. would like to acknowledge useful discussions with Ulrich Zülicke at the early stages of this work. M. Y. and J. K. F. were supported by the Department of Energy, Office of Basic Energy Sciences, Division of Materials Sciences and Engineering under Contract No. DE-FG02-08ER46542 for the formal developments, the numerical work, and the writing of the manuscript. J. K. F. was also supported by the McDevitt bequest at Georgetown University. J. L. was supported by the Research Council of Norway through its Centres of Excellence funding scheme Grant No. 262633 and Grant No. 353894.